a) Discuss why newborns homozygous for β-thalassemia are often healthy, with severe disease symptoms developing during the first several months after
birth, while homozygotes for α-thalassemia are affected from birth.

b) A couple suffer the prenatal loss of their baby due to Barts hydrops fetalis.
Describe the molecular basis of this condition. Draw a pedigree with genotypes
that illustrates to the carrier parents the genetic basis of the infant’s
thalassemia, and the likely outcomes for any future pregnancies. Explain why
a Melanesian couple, who both also have the α-thalassemia trait, are unlikely
to have a similarly affected infant
Answer #1

Answer :-

(a) New borns homozygoous for \beta thalassemia are often healthy, with severe disease symptoms developing during the first several months after birth because at the birth baby seems to be normal with major thalassemia. The reason being is that at birth fetal Hemoglobin is present which having two alpha chains and two gamma chains. As there is absence of beta chain that protects the baby at birth from the effect of thalassemia major. Effect of thalassemia begins to occur during the first several months after birth. During that time, newborn baby unable to grow properly, intestinal problems,problem in feeding due to lack of oxygen from the thalassemia effect to the newborn baby whereas homozygotes for a \alpha thalassemia are affected from birth and the reason being is that four missing or mutated gene leads to the condition of \alpha thalassemia major and it is also called hydrops fetails. This condition leads to the death of fetus before delivery or new born baby will die after some time. In this condition body of New born baby has excess fluid.

(b) Barts hydrops fetalis is caused by the inactivation of four-alpha globin alleles which causes deficiency in alpha globin chains of hemoglobin. Deficiency in Hemoglobin tetramers also takes place which leads to incapable of oxygen delivery in tissues. Mutation in HBA1 and HBA2 genes occur. In this condition, the death generally occurs after birth or in uterus. Those new born which survive show abnormal development and limb defects.

Pedigree can't be drawn as information is not properly given.

\alpha genotype being common in many population including Melanesian couple. The parent with this genotype can't transmit it to its offspring. In homozygotes \alpha thalassemia, the functional gene is 0, so the production of \alpha chain will be 0%.